OpenStudy (anonymous):

I dont want the answer, I just want to know how to find it. Solve for x: −3(x + 3) = −3(x + 1) − 5 A. 6 B. −6 C. All real numbers D. No solution Also, what does it mean by all real numbers? I didnt see that in the lesson.

OpenStudy (anonymous):

Do you know the distributive property?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Distribute where possible to get -3x - 9 = -3x - 3 -5 After that I'm sure you've got it.

OpenStudy (anonymous):

@gabriel11 openstudy.com/code-of-conduct and that's not right

OpenStudy (anonymous):

do you still want to know how to work the question out, even though this gabriel kid can't follow rules?

OpenStudy (anonymous):

Please.. I did the distributive property, but like I said, im not sure what it means by all real numbers soo

OpenStudy (anonymous):

well, a real number is any irrational or rational number.

OpenStudy (anonymous):

Irrational number = a real number that cannot be expressed as a rational number ex. sqrt of 2 or PI Rational number = an integer or a fraction. 3/2 or 7

OpenStudy (anonymous):

Okay, that makes sence(: Soo, once I get -3x + -9 = -3x + -3 - 5, what would I do next? Thats where I get lost.

OpenStudy (anonymous):

Personally i would STOP and analyze that question, and realize the answer right there but we combine like terms -3x - 9 = -3x - 8 add 8 to each side -3x - 1 = -3x then add 3x to each side -1 = 0

OpenStudy (anonymous):

why i said we should analyze the question is because we can already see that the answer will be equal to 0 because the -3x's will cancel each other.

OpenStudy (anonymous):

I see that now, thank you soo much